A 1-kg block of ice is at -15° C. 4 x 105 J of heat is added to the ice block. NOTE: (a) should be divided into two parts: how much heat is needed to raise the temperature of the ice to 0 °C, and how much heat is required to melt it; then compare that sum total to 4.0 x 105 J (the specific heat to use for ice is provided in the accompanying sheet to this homework)

(a) Does it fully melt?
(b) What is the final overall temperature?
1
Expert's answer
2020-04-28T09:48:32-0400

a)


"Q_1=(1)(2060)(15)=30900\ J"

"Q_2=(1)(335000)=335000\ J"

"Q=335000+30900=365900\ J<400000\ J"

So, all ice melted.

b)


"(1)(4200)(T-0)=400000-365900"

"T=8.1\degree C"